Grade 12th passMechanics

The resultant of two forces acting at a point is 65 kN. It is observed that one force is double than that of the other and if the direction of one of the is reversed the resultant becomes 45kN. Find the magnitude of the forces and angle between them ?

Chris George
6 Years agoGrade 12th pass

1 Answer

Vikas TU
6 Years ago
Dear student 
Let the forces be of magnitude x and 2x. Then, 
From Prallelogram law 
x^2 +4x^2+4x^2cosy=(65)^2
and similarly 
x^2+4x^2-4x^2cosy=(45)^2
Simplifying we get,
10x2=6250 or x=25
and 
8x2cosy=2200
or cosy=44/100
so forces are 25N and 50N and at an angle of somewhere round 64 degrees.